# Using Fiber Optic Bundles to Miniaturize Vision-Based Tactile Sensors

### v2

@@ Abstract
an optical system design that integrates fibers for imaging and illumination and a distal lens to achieve a wide field of view in a hemispherical dome;
@@ II-A Vision-Based Tactile Sensors
To better suit dexterous manipulation tasks, recent works have developed these sensing designs from planar to a three-dimensional. One common form factor that has emerged is an elastomer gel shaped in a hemisphere or dome, reminiscent of a human fingertip that may then be mounted on a robotic finger. OmniTact [6], GelTip [12], and GelSight360 [7] are all recent examples of these fingertip vision-based tactile sensors. Another common theme is that of a slim elastomer gel for direct integration with a gripper, with elastomers shaped as wedges [4], finrays [13], or fingers themselves [8, 14].
@@ II-B Fiber-Based Tactile Sensing
In this work, we revisit the idea of fiber-based tactile sensing through the lens of endoscopic design principles to achieve high spatial resolution with small sensor size. Modern incarnations of flexible fibreoptic endoscopes often employ a proximal video camera for image capture through the CFB [17], or use a miniature CCD chip at the distal end of a flexible shaft, with incoherent optical fiber bundles to deliver diffuse white-light illumination from the proximal to distal end [26]. Both designs result in a slim imaging scope, with trade-offs between size and resolution. While previous research in fiber-based imaging has investigated optical techniques (e.g., imaging relay and scanning mechanisms [27], fiber characterization [28], and case studies in medicine [29, 30]), we do not know of works that extend these design details to the system level needed for integration into vision-based tactile sensors. Such details include the arrangement of fibers and lighting for a hemispherical elastomer dome and the design of a miniature distal lens for imaging at close range with high field of view.
@@ II-C Tactile Applications in Medicine
Nevertheless, because the DRE and similar palpation examinations are widely accepted as subjective measurements, researchers have investigated ways to quantify them with sensorized robotic systems. To mimic the DRE, some researchers have affixed a probe to a force-torque sensor for palpation case studies, or have used fibers with Fiber Bragg Gratings (FBGs) for strain sensing [42, 43, 44]. Others have investigated robotic palpation classifiers for breast and prostate tumors [45, 46]. None of these investigations has demonstrated a high-resolution vision-based tactile sensor small enough to be clinically relevant for the DRE and similarly constrained manipulation tasks.
@@ III Design and Fabrication
In this section we detail the design principles and corresponding choices made during the design of the DIGIT Pinki. We additionally describe how each of the components of the sensor works, and present their manufacturing process.
@@ III-C Illumination
Consideration was given to the location of the illumination fibers relative to the base of the elastomer gel dome. We defined their placement to the elastomer gel with these parameters: number of fiber bundles , radius of each fiber bundle (inclusive of any jackets) , the radial distance between the center of each bundle from sensor center , and angular spacing between bundles . We empirically found that while 3 was a minimal number of bundles (for RGB illumination), 6 fiber bundles equally spaced at radius worked better.
@@
One result from using fiber bundles for lighting is that the output light profile is different from a Lambertian diffuse source. To study this, we modeled in Zemax Optics Studio a comparison between the 5050 LED package without fibers and with the diameter PMMA fiber bundle as a lighting conduit (Fig. 3). In this model, we plotted the ray-traced light intensity received at a rectangular detector away from the light source using 1,000 analysis rays for both scenarios. We found that illumination through a fiber bundle is more narrow in output compared to the Lambertian light source of a direct LED. We then approximated the lighting as a Gaussian beam instead of a Lambertian source, with a cone of light given by the numerical aperture () of the fiber bundle. Less uniform illumination is not necessarily disadvantageous; while hotspots may saturate the image sensor more quickly, reducing the maximum detectable deformation, they can also increase imaging contrast for perceiving fine features.
@@ III-D Imaging
We chose the IMX298 16MP Arducam camera with a pixel size of 1.12 × and a maximum image resolution of 4656 × 3496 pixels when imaging at , though the frame rate is adjustable up to at lower resolutions. The camera parameters such as brightness, white balance, and exposure are also adjustable.
For imaging through a fiber bundle, the effective spatial resolution is most often constrained by the fiber packing density and fiber core diameter, rather than the camera sensor and imaging optics. Resolution in this case refers to the minimum feature dimensions that can be resolved with a modular transfer function (MTF) 0.5. In the proof of concept, we used a 7400-core imaging fiber bundle made of Poly (methyl methacrylate) plastic, which was packed as a hexagonal array in a diameter polyethylene jacket and had a individual fiber core of .
To achieve a high field of view of the hemispherical gel, a wide-angle lens is necessary at the distal end of the fiber. Most existing vision-based tactile sensors use a commercial off-the-shelf camera module with a M12 wide angle lens, but these lenses are too wide for our purposes.
Thus, we designed a custom compact hyperfisheye lens for the DIGIT Pinki, as shown in Fig. 4, which provided a field of view. This custom lens was optimized for tactile sensing input in a dome-shaped sensor. In comparison to commercially-available lenses intended for photography, we were relatively unconcerned about optical distortions. Therefore the lens intentionally allowed chromatic aberrations and removed anti-reflective coatings to capture reflection and scattering inside the elastomer. The design used APEL 5014 plastic material and could be manufactured either by diamond turning or a molding process.
Due to resource constraints, we were unable to manufacture the hyperfisheye lens design in time for testing. Therefore, for tests described in the next section, we only image the tip ( field of view) with a non-hyperfisheye distal lens. The spatial resolution result reported in Section IV is with the lens, but the same resolution with the hyperfisheye lens could be attained by increasing the fiber core count accordingly.
@@ III-E Assembly
We assembled our prototype, as pictured in Fig. 2, in the following manner.
@@
Next, we prepared the proximal end of the sensor, consisting of the proximal illumination and imaging systems. The illumination and imaging systems may be co-located, but in the prototype built in this work they were separated for modularity.
@@
To image the proximal end face of the imaging fiber bundle, we set up an optical bench system with ThorLabs components. All pieces were mounted to an optical breadboard, which we set up with a adjustable optical cage and light shields. From the imaging fiber bundle, we aligned an adjustable diopter, a 10x Plan microscope objective, and a microscope eyepiece to a 16MP IMX298 USB camera connected to a computer.
@@ IV-A Spatial Resolution
[…]
| DIGIT |
304 |
60 |
0.150 |
0.006 |
0.012 |
| DIGIT Pinki |
1,404* |
10** |
0.22 |
0.005 |
0.005 |
This number is from the proposed hyperfisheye lens design.
Sampling rate at highest resolution of 4656×3496 pixels. A sampling rate is available at lower resolutions (1920×1080 pixels).
To verify the resolution of the sensor system optics, we used a standard United States Air Force (USAF) resolution target as commonly used for microscopes and cameras. This benchtop test is a tri-bar test, where the optical resolution of the system is determined by assessing the visibility of groups of bars. The resolution is the smallest group where all three black bars can be separated. The image from the USAF resolution target test captured through the fiber bundle is shown in Fig. 5a with a working distance equal to that at the tip of the gel. We clearly observed Group Number 1, Element Number 2, for a resolution of 2.24 lines per . Based on this resolution test target, the DIGIT Pinki can resolve a line width of approximately . The true resolution limit is between and (which corresponds to Element Number 3). Note that this resolution limit is for a system using the distal lens and 7,400-core imaging fiber; changes to either optical components could improve the resolution limit.
We also tested the resolution of imaging through an elastomer gel. To do so, we custom machined a calibration test target block. The block has line widths of , , , and . The results from this test are shown in Fig. 5b-e. The camera exposure in this sequence is set to . The indentations of the lines are clearly visible down to .
@@ IV-B Normal and Shear Force Estimation
Normal |
Flat |
Normal (10k) |
Flat |
Shear |
Flat |
Shear (10k) |
@@
We used an indentor to probe DIGIT Pinki’s sensing element. As shown in Fig. 5(a), we mounted the DIGIT Pinki on a Mecha robotic arm, and installed the metal indenter probe on a force sensor on top of a linear stage. During this process of contacting the surface, we collected images of the surface under contact and the contact force vectors from a force sensor embedded under the indenter. We used three different indenter probe tips: diameter, (on the order of the size of the sensor tip), and a flat metal surface (larger than the sensor tip). The variety in tip sizes produces different contact loading conditions.
We collected two types of datasets: normal and shear.
@@ IV-B2 Shear Force Dataset
For shear forces, we used a three-axis force sensor to collect the contact force vector. The procedure is similar to normal force collection at first, with an additional step of loading the probe in shear after normal force is loaded. First, the probe was controlled to apply normal force, and then it was moved tangentially to the surface loading shear force up to . Finally, the probe was moved back to the previous location, unloading the shear force. If there was non-zero residual shear force, we discarded the data because slip might have occurred. In total, we collected 138,348 image-force pairs over the course of several days to create our shear dataset. We randomly split this into a training set (100k samples), validation set (10k samples), and test set (remaining 28k samples). We also prepared a smaller subset of 10,000 randomly-selected image-force pairs with the same split, which we refer to as “Shear (10k).”
@@
We report the normal and shear force resolution in Table I, and the RMSE values in Table II. We also report a force-angle precision for the full shear dataset, and a force-angle precision for the smaller 10k-sample dataset. We report a median absolute error of for the normal over unseen test sessions, and for the shear over unseen test sessions.
Based on our results, DIGIT Pinki is able to accurately estimate light normal and shear forces over the tip of the elastomer surface. There is no significant difference in performance between the different indenter probe sizes, suggesting that DIGIT Pinki would perform well over a variety of contact sizes. The model is also able to learn comparably from both the full dataset and the smaller dataset (10k samples) for this task. In the next subsection, we also investigate an analytic model that supports the accurate force estimation performance in the light force regime.
@@ IV-C Contact Area Model
Based on our results, the Hertzian contact model trend fits closely to the observed contact area radius trend for light forces. This suggests that the gel elastic behavior is approximately linear while operating in the light force regime, which aids in learning or analytically solving for contact conditions. We note, however, that this sensor, like other rounded optical tactile sensors, will experience a reduction in sensitivity with higher loads at which the gel deformation becomes less linear and the contact patch grows more slowly. Accordingly it may be desirable to tune the sensor tip geometry to the expected geometry of an indenting object and to the anticipated load range.
@@ V Silicone Palpation Experiments
For a classification task with our sensor, we desire the ability to classify firmness over a range of values that represent both healthy and unhealthy tissue. For this, we prepared a custom dataset associating stiffness (defined as hardness values on the Shore durometer scale for elastomers) to silicone samples of different relevant geometries. Using this dataset, we trained a classification model that classifies the hardness values based on a sequence of 16 image frames, in order to show DIGIT Pinki’s potential for palpation use cases.
@@ V-A Hardness Dataset Collection
To prepare the silicone hardness dataset, we cast custom silicone samples with varied hardness values and geometries by mixing different quantities of Smooth-on Ecoflex 0050 and Smooth-On Smooth-Sil 945 platinum silicones. Ecoflex 0050 (50 Shore 00) was chosen to represent what would be considered healthy tissue, since it is reported to have a similar elastic modulus to real healthy prostate tissues from patients [50]. Smooth-Sil 945 silicone was chosen because it is significantly harder, has a fast cure time, and has a 1:1 mixing ratio that makes it easy to custom-blend different ratios. The exact mixing ratios used are detailed in the Appendix C.
We prepared molds for four different types of surfaces: flat, and with bumps of nd diameter, all raised above the surface. The bump sizes were chosen based on the different diameter categories reported for lesions in prostate tissues: we desired one bump under , one between 5 and , and one over . The prostate posterior is typically smooth and flat or gently concave, which the flat surface sample mimics. The flat surface sample is also used in a durometer test to find an exact hardness value for each batch.
The molds were printed with a FormLabs 3 printer using Rigid 1K resin and prepared with XTC-3D brush-on coating. We then poured the silicone mixtures into the four different molds to produce four silicone samples per batch. The samples are thick and 25 × 25 across; examples are shown in Fig. 7(a).
In total, 11 sets of silicone samples (44 individual swatches) were prepared in this manner, reflecting 11 different hardness values. The samples were randomly colored with Silc-Pig pigment (0.5% by weight) to help distinguish between them, but these addition of these colorants did not affect the hardness value. We assigned the hardness values to each sample batch by taking the average of five measurements with the durometer on the flat surface sample, where the 0.5A sample is the healthy tissue baseline.
We then used the experimental setup shown in Fig. 7(b) to manually palpate the silicone samples for the dataset. As shown in the figure, the samples were mounted on a 3D-printed holder that was rigidly bolted to an ATI Gamma force-torque sensor. During the palpation session, we used a computer to synchronously record both image and force data.
The palpation technique we used for the dataset collection was to manually make and break contact with the sensor tip normal to the surface of the silicone feature. We gently palpated (no more than about normal force) each silicone sample feature three times for one minute each, mimicking the gentleness necessary for patient comfort in a real examination. Each palpation session consisted of the raw image sequence from the tactile sensor and the forces from the ATI Gamma. We did not use the force data during model training.

### v3

@@ Abstract
an optical system design that integrates fibers as conduits for imaging and illumination to separate the sensing element from supporting circuitry;
@@ II-A Vision-Based Tactile Sensors
To better suit dexterous manipulation tasks, recent works have developed these sensing designs from planar to a three-dimensional. One common form factor that has emerged is an elastomer gel shaped in a hemisphere or dome, reminiscent of a human fingertip that may then be mounted on a robotic finger. OmniTact [6], GelTip [12], and GelSight360 [7] are all recent examples of these fingertip vision-based tactile sensors with an internal rigid shell that provides additional structural support. A hemispherical shape may also be achieved without an internal rigid shell, of which Soft-Bubble [13] and DenseTact [14] are examples. Another common theme is that of a slim elastomer gel for direct integration with a gripper, with elastomers shaped as wedges [4], finrays [15], or fingers themselves [8, 16].
@@ II-B Fiber-Based Tactile Sensing
In this work, we revisit the idea of fiber-based tactile sensing through the lens of endoscopic design principles to achieve high spatial resolution with small sensor size. Modern incarnations of flexible fiberoptic endoscopes often employ a proximal video camera for image capture through the CFB [19], or use a miniature CCD chip at the distal end of a flexible shaft, with incoherent optical fiber bundles to deliver diffuse white-light illumination from the proximal to distal end [28]. Both designs result in a slim imaging scope, with trade-offs between size and resolution. While previous research in fiber-based imaging has investigated optical techniques (e.g., imaging relay and scanning mechanisms [29], fiber characterization [30], and case studies in medicine [31, 32]), we do not know of works that extend these design details to the system level needed for integration into vision-based tactile sensors. Such details include the arrangement of fibers and lighting for a hemispherical elastomer dome and the design of a miniature distal lens for imaging at close range with high field of view.
@@ II-C Tactile Applications in Medicine
Nevertheless, because the DRE and similar palpation examinations are widely accepted as subjective measurements, researchers have investigated ways to quantify them with sensorized robotic systems. To mimic the DRE, some researchers have affixed a probe to a force-torque sensor for palpation case studies, or have used fibers with Fiber Bragg Gratings (FBGs) for strain sensing [44, 45, 46]. Although these efforts are promising, clinicians today practice a variety of palpation techniques involving both normal and shear forces to obtain a wealth of information about the tissue (e.g. firmness, texture, temperature), which a simple 1-DOF force sensor may not be able to capture. Other researchers have investigated robotic palpation classifiers for breast and prostate tumors [47, 48]. None of these investigations has demonstrated a high-resolution vision-based tactile sensor small enough to be clinically relevant for the DRE and similarly constrained manipulation tasks.
@@ III Design and Fabrication
This section details the design principles and corresponding choices made during the design of the DIGIT Pinki. Descriptions of how each of the components of the sensor works and their manufacturing process are presented.
@@ III-C Illumination
Consideration was given to the location of the illumination fibers relative to the base of the elastomer gel dome. We defined their placement to the elastomer gel with these parameters: number of fiber bundles , radius of each fiber bundle (inclusive of any jackets) , the radial distance between the center of each bundle from sensor center , and angular spacing between bundles . While 3 was a minimal number of bundles for RGB illumination (the color channels provide shadow and contrast to aid models to learn from an image), the following parameters empirically worked to provide illumination coverage across the field of view: , , , and .
@@
One result from using fiber bundles for lighting is that the output light profile is different from a Lambertian diffuse source. To study this, we modeled in Zemax Optics Studio a comparison between the 5050 LED package without fibers and with the diameter PMMA fiber bundle as a lighting conduit (Fig. 3). In this model, we plotted the ray-traced light intensity received at a rectangular detector away from the light source using 1,000 analysis rays for both scenarios. We found that illumination through a fiber bundle is more narrow in output compared to the Lambertian light source of a direct LED. We then approximated the lighting as a Gaussian beam instead of a Lambertian source, with a cone of light given by the numerical aperture () of the fiber bundle. The amount of Gaussian scatter affects the image contrast and therefore sensitivity of the sensor. Very low amounts of Gaussian scatter may result in non-uniform background illumination and specular glints on the sides of the sensor that could saturate the vision system, which is not necessarily disadvantageous. However, to achieve a high contrast across the field of view of the sensor, some degree of Gaussian scatter is needed for imaging contrast across the field of view. Accurate scattering can be realized with proper tool machining; sandblasting the mold is a more accessible method with adequate results.
@@ III-D Imaging
We chose the IMX298 16MP Arducam camera with a pixel size of 1.12 × and a maximum image resolution of 4656 × 3496 pixels when imaging at , though the frame rate is adjustable up to at lower resolutions. The camera parameters such as brightness, white balance, and exposure are also adjustable. At 38 x , this particular camera would be too bulky to be deployed at the base of the gel dome in the style of the more typical vision-based tactile sensors.
For imaging through a fiber bundle, the effective spatial resolution is most often constrained by the fiber packing density and fiber core diameter, rather than the camera sensor and imaging optics. Resolution in this case refers to the minimum feature dimensions that can be resolved with a modulation transfer function (MTF) 0.5. The proof of concept used a 7400-core imaging fiber bundle made of Poly (methyl methacrylate) plastic, which was packed as a hexagonal array in a diameter polyethylene jacket and had a individual fiber core of .
To achieve a high field of view of the hemispherical gel, a wide-angle lens is necessary at the distal end of the fiber. Most existing vision-based tactile sensors use a commercial off-the-shelf camera module with a M12 wide angle lens, but these lenses are too bulky for our purposes.
For tests described in the next section, we only image the tip ( field of view) with a non-hyperfisheye distal lens. However, it is possible to substitute in a hyperfisheye lens, assuming it could be obtained at moderate cost, which we leave for future iterations of this work.
@@ III-E Assembly
The prototype, as pictured in Fig. 2, was assembled in the following manner.
@@
Next, we prepared the proximal end of the sensor, consisting of the proximal illumination and imaging systems. The illumination and imaging systems may be co-located, but in the prototype built in this work they were separated for easier access to the individual components for debugging.
@@
To image the proximal end face of the imaging fiber bundle, we set up an optical bench system with ThorLabs components. All pieces were mounted to an optical breadboard, which we set up with a adjustable optical cage and light shields. From the imaging fiber bundle, we aligned an adjustable diopter, a 10x Plan microscope objective, and a microscope eyepiece to a 16MP Arducam IMX298 USB camera connected to a computer. The lens of the Arducam IMX298 USB camera was focused manually before assembly. The adjustable diopter was used for fine focusing of the output image, and the placement of the microscope objective and eyepiece on the optical breadboard was also adjusted to focus the output image.
@@ IV-A Spatial Resolution
| DIGIT | 304 | 60 | 0.150 | 0.006 | 0.012 |
| DIGIT Pinki | 181* | 10* | 0.22 | 0.005 | 0.005 |
Calculation with distal lens.
Sampling rate at highest resolution of 4656×3496 pixels. A sampling rate is available at lower resolutions (1920×1080 pixels), which is the frequency and resolution used for data collection in this work.
To verify the resolution of the sensor system optics, we used a standard United States Air Force (USAF) resolution target as commonly used for microscopes and cameras. This benchtop test is a tri-bar test, where the optical resolution of the system is determined by assessing the visibility of groups of bars. The resolution is the smallest group where all three black bars can be separated. The image from the USAF resolution target test captured through the fiber bundle is shown in Fig. 4a with a working distance equal to that at the tip of the gel. We clearly observed Group Number 1, Element Number 2, for a resolution of 2.24 lines per . Based on this resolution test target, the DIGIT Pinki can resolve a line width of approximately . The true resolution limit is between and (which corresponds to Element Number 3). Note that this resolution limit is for a system using the distal lens and 7,400-core imaging fiber; changes to either optical component could improve the resolution limit.
We also tested the resolution of imaging through an elastomer gel. To do so, we custom-machined a calibration test target block. The block has line widths of , , , and . The results from this test are shown in Fig. 4b-e. The camera exposure in this sequence is set to . The indentations of the lines are clearly visible down to .
@@ IV-B Normal and Shear Force Estimation
| Normal | ||
| Flat | ||
| Normal (10k) | ||
| Flat | ||
| Shear | ||
| Flat | ||
| Shear (10k) | ||
@@
We used an indentor to probe DIGIT Pinki’s sensing element. As shown in Fig. 5(a), we mounted the DIGIT Pinki on a six-axis industrial robot arm (Meca500) with a precision of . The metal indenter probe was installed atop the force sensor such that the sensor frame and probe frame are axis aligned,and the force sensor was tared to the no-contact condition. This assembly was mounted on a hexapod 6-axis stage (Newport HXP50) which can be precisely controlled to translate with and rotate with increments, but during data collection the stage was stationary and the robot arm translated. More details on the robot indenter data collection setup can be found in Appendix D. During this process of contacting the sensor surface, we collected images of the surface under contact and the contact force vectors from the force sensor embedded under the indenter. We used three different indenter probe tips: diameter, (on the order of the size of the sensor tip), and a flat metal surface (larger than the sensor tip). The variety in tip sizes produces different contact loading conditions.
Two types of datasets were collected: normal and shear.
@@ IV-B2 Shear Force Dataset
For shear forces, we used a three-axis force sensor (FUTEK QMA147 3-axis sensor) to collect the contact force vector. The procedure is similar to normal force collection at first, with an additional step of loading the probe in shear after normal force is loaded. First, the probe was controlled to apply normal force, and then it was moved tangentially to the surface loading shear force up to . Finally, the probe was moved back to the previous location, unloading the shear force. If there was non-zero residual shear force, we discarded the data because slip might have occurred. In total, we collected 138,348 image-force pairs over the course of several days to create our shear dataset. We randomly split this into a training set (100k samples), validation set (10k samples), and test set (remaining 28k samples). We also prepared a smaller subset of 10,000 randomly-selected image-force pairs with the same split, which we refer to as “Shear (10k).”
@@
The normal and shear force resolution are reported in Table I, and the RMSE values in Table II. We also report a force-angle precision for the full shear dataset, and a force-angle precision for the smaller 10k-sample dataset. We report a median absolute error of for normal force, and for shear force based on the random train/validation/test split.
An additional study was conducted on the 10k dataset, where instead of randomly splitting the dataset into the train/validation/test sets, we explicitly use image-force pairs from one spatial point as the validation set and image-force pairs from another spatial point as the test set. This guarantees that the training, validation, and test sets are all from different spatial locations, which are detailed in Appendix F. The model achieved a median absolute error of , a comparable performance with the randomized split.
Based on our results, DIGIT Pinki is able to accurately estimate light normal and shear forces over the tip of the elastomer surface. There is no significant difference in performance between the different indenter probe sizes, suggesting that DIGIT Pinki would perform well over a variety of contact sizes. The model is also able to learn comparably from both the full dataset and the smaller dataset (10k samples) for this task, and can learn to predict force on an unseen spatial location at the tip of the elastomer surface. In the next subsection, we also investigate an analytic model that supports the accurate force estimation performance in the light force regime.
@@ IV-C Contact Area Model
Based on our results, the Hertzian contact model trend fits closely to the observed contact area radius trend for light forces. This suggests that the gel elastic behavior is approximately linear while operating in the light force regime, which aids in learning or analytically solving for contact conditions. We note, however, that this sensor, like other rounded optical tactile sensors, will experience a reduction in sensitivity with higher loads at which the gel deformation becomes less linear and the contact patch grows more slowly. Accordingly, it may be desirable to tune the sensor tip geometry to the expected geometry of an indenting object and to the anticipated load range.
@@ V Silicone Palpation Experiments
For a classification task with our sensor, we desire the ability to classify firmness over a range of values that represent both healthy and unhealthy tissue. For this, we prepared a custom dataset associating stiffness (defined as hardness values on the Shore durometer scale for elastomers) to silicone samples of different relevant geometries. Using this dataset, a classification model was trained that classifies the hardness values based on a sequence of 16 image frames, in order to show DIGIT Pinki’s potential for palpation use cases. An ablative analysis is also provided on the effect of the contact patch and image sequence length to performance.
@@ V-A Hardness Dataset Collection
To prepare the silicone hardness dataset, custom silicone samples were casted with varied hardness values and geometries by mixing different quantities of Smooth-on Ecoflex 0050 and Smooth-On Smooth-Sil 945 platinum silicones. Ecoflex 0050 (50 Shore 00) was chosen to represent what would be considered healthy tissue, since it is reported to have a similar elastic modulus to real healthy prostate tissues from patients [52]. Smooth-Sil 945 silicone was chosen because it is significantly harder, has a fast cure time, and has a 1:1 mixing ratio that makes it easy to custom-blend different ratios. The exact mixing ratios used are detailed in the Appendix E.
Molds were prepared for four different types of surfaces: flat, and with bumps of nd diameter, all raised above the surface. The bump sizes were chosen based on the different diameter categories reported for lesions in prostate tissues: we desired one bump under , one between 5 and , and one over . The prostate posterior is typically smooth and flat or gently concave, which the flat surface sample mimics. The flat surface sample is also used in a durometer test to find an exact hardness value for each batch.
The molds were printed with a FormLabs 3 printer using Rigid 1K resin and prepared with XTC-3D brush-on coating. The silicone mixtures were then poured into the four different molds to produce four silicone samples per batch. The samples are thick and 25 × 25 across; examples are shown in Fig. 7(a).
In total, 11 sets of silicone samples (44 individual swatches) were prepared in this manner, reflecting 11 different hardness values. The samples were randomly colored with Silc-Pig pigment (0.5% by weight) to help distinguish between them, but these addition of these colorants did not affect the hardness value. The hardness values were assigned to each sample batch by taking the average of five measurements with the durometer on the flat surface sample, where the 0.5A sample is the healthy tissue baseline.
The experimental setup shown in Fig. 7(b) was used to manually palpate the silicone samples for the dataset. As shown in the figure, the samples were mounted on a 3D-printed holder that was rigidly bolted to an ATI Gamma force-torque sensor. During the palpation session, a computer synchronously recorded both image and force data.
The palpation technique we used for the dataset collection was to manually make and break contact with the sensor tip normal to the surface of the silicone feature. We gently palpated (no more than about normal force) each silicone sample feature three times for one minute each, mimicking the gentleness necessary for patient comfort in a real examination. Each palpation session consisted of the raw image sequence from the tactile sensor and the forces from the ATI Gamma. The force data was not used during model training.